8. How much more is an annuity of $12,000 annually for 10 years with an interest of 6% than a perpetuity of $8,000 at the same rate?
Answer #1

Annuity =12000
Number of years =10
Rate =6%
PV of annuity =PMT*((1-(1+r)^-n)/r)=12000*((1-(1+6%)^-10)/6%)=88,321.04

PV of Perpetuity =Perpetuity/Rate =8000/6% =133,333.33